If a 90% confidence interval for the mean is found to be (1.25,4.75) using a sample of n=100, find the 95% confidence interval for the mean. a. Is this interval narrower or wider? Why? b. In general, explain how the sample size and the confidence level affect the width of a confidence interval. 


1
Expert's answer
2021-06-09T15:23:07-0400

Assume that the population standard deviation is known.

"CI=\bar X\pm E"

for the 90% CI, "E=\frac{4.75-1.25}{2}=1.75"

And "\bar X=1.25+1.75=3"

But "E=z_{\frac{\alpha}{2}}\frac{\sigma}{\sqrt{n}}"

For 90% confidence level, "z=1.645"

Thus,"1.75=1.645(\frac{\sigma}{\sqrt{100}})" solving for sigma,

"\sigma=10.64"

a. 95% CI

For 95% CI

"Z_{\frac{\alpha}{2}}=1.96"

"E=1.96(\frac{10.64}{\sqrt{100}})=2.085"

"CI=3\pm2.085=(0.915,5.085)"

95% CI is wider than 90% CI.

b.

Increasing sample size decreases the width of the confidence interval since as sample size increases, sample mean approaches the population mean. however, increasing confidence level widens the width of the confidence interval since high confidence occurs when one is less precise.